three carts of masses m1+4.00 kg, m2+10.0 kg and m3 = 3.00 kg move on a frictionless, horizontal track with speeds of v1=5.00 m/s to the right, v2 =3.00 m/s to the right, and v3= 4.00 m/s to the left. Velcro couples make the carts stick together after colliding. Find the final velocity of the train of three carts. (b) What if? does your answer in part (a) require that all the carts collide and stick


1
Expert's answer
2021-05-31T15:38:15-0400

Solution.

"m_1=4.00kg;"

"m_2=10.0kg;"

"m_3=3.00kg;"

"v_1=5.00m\/s;"

"v_2=3.00m\/s;"

"v_3=-4.00m\/s;"

"m_1v_1+m_2v_2+m_3v_3=(m_1+m_2+m_3)v\implies"

"v=\dfrac{m_1v_1+m_2v_2+m_3v_3}{m_1+m_2+m_3};"

"v=\dfrac{4.00kg\sdot5.00m/s+10.0kg\sdot3.00m/s-3.00kg\sdot4.00m/s}{4.00kg+10.0kg+3.00kg}=2.24m/s;"

Answer:"2.24m\/s."